Intrinsic Value of Merger target

Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1.5 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.25 (given its target capital structure). Vandell has $8.80 million in debt that trades at par and pays a 7% interest rate. Vandell’s free cash flow (FCF0) is $2 million per year and is expected to grow at a constant rate of 4% a year. Vandell pays a 25% combined federal-plus-state tax rate, the same rate paid by Hastings. The risk-free rate of interest is 4%, and the market risk premium is 8%. Hasting’s first step is to estimate the current intrinsic value of Vandell.

  1. What is Vandell’s cost of equity? Do not round intermediate calculations. Round your answer to two decimal places.

      %

  2. What is its weighted average cost of capital? Do not round intermediate calculations. Round your answer to two decimal places.

      %

  3. What is Vandell’s intrinsic value of operations? (Hint: Use the free cash flow corporate valuation model.) Enter your answer in millions. For example, an answer of $1.23 million should be entered as 1.23, not 1,230,000. Do not round intermediate calculations. Round your answer to two decimal places.

    $   million

  4. Based on this analysis, what is the minimum stock price that Vandell’s shareholders should accept? Do not round intermediate calculations. Round your answer to the nearest cent.

    $   / share
Answer #1

a. As per Capital Asset Pricing Model (CAPM):
Cost of Equity (Ke) = Risk Free Rate (Rf) + [Beta * Market Risk Premium]
= 0.04 + (1.25 * 0.08)
= 0.14
Therefore, Cost of Equity (Ke) = 0.14 or 14%

b. Since the debt is trading at part, the pre-tax cost of debt will same as the rate of Interest.
Therefore, Cost of Debt (Kd) = 7% (Pre-tax)
Post Tax Cost of Debt (Kd) = 7*(1 - tax rate)
= 7*(1-0.25)
Post tax cost of debt (Kd) = 5.25%

Vandell's target structuring consists of 30% debt, hence it conists of (100-30) 70% of equity.

Therefore,
Vandell's WACC = (Kd*30%) + (Ke*70%)
= (5.25 * 30%) + (14 * 70%)
= 11.375 or 11.38%

Therefore, Vandell's Weighted Average Cost of Capital is 11.38%

c. Vandell's Free Cash flows are $2m per year and the growth rate is expected to be constant at 4%
Therefore,
Intrinsic Value of Operations = (Cash Flow of Year0 + Growth Rate) / (WACC - Growth Rate)
= ($2m + 2m*4%) / (11.38% - 4%)
= $28.18m

Therefore, Vandell's Intrinsic Value of Operations is $28.18 million.

d. Intrinsic Value of Vandell's business = $28.18million
Value of Debt = $8.80 million
Therefore, Value of Equity = Value of Business - Value of Debt
= $28.18m - $8.80m
= $19.38m

Total Number of shares outstanding = 1.5 million
Therefore, Value per share = $19.38m / 1.5m
= $12.92 or $12.9

Therefore, the minimum stock price that Vandell's shareholders should accept is $12.9 per share